### v1

# Examining Multilingual Embedding Models Cross-Lingually Through LLM-Generated Adversarial Examples

Abstract
The evaluation of cross-lingual semantic search capabilities of models is often limited to existing datasets from tasks such as information retrieval and semantic textual similarity. To allow for domain-specific evaluation, we introduce Cross Lingual Semantic Discrimination (CLSD), a novel cross-lingual semantic search task that requires only a set of parallel sentence pairs of the language pair of interest within the target domain. This task focuses on the ability of a model to cross-lingually rank the true parallel sentence higher than hard negatives generated by a large language model. We create four instances of our introduced CLSD task for the language pair German-French within the domain of news. Within this case study, we find that models that are also fine-tuned for retrieval tasks (e.g., multilingual E5) benefit from using English as the pivot language, while bitext mining models such as LaBSE perform best directly cross-lingually. We also show a fine-grained similarity analysis enabled by our distractor generation strategy, indicating that different embedding models are sensitive to different types of perturbations.
Examining Multilingual Embedding Models Cross-Lingually Through LLM-Generated Adversarial Examples
Andrianos Michail, Simon Clematide, Rico Sennrich University of Zurich {andrianos.michail,simon.clematide,sennrich}@cl.uzh.ch
1 Introduction
When considering which model to use to perform semantic search within a specific domain and language pair, standard benchmarks such as MTEB Muennighoff et al. (2023) do not cover every domain within all language pairs. We posit that one should aim to directly evaluate semantic search for the target language pair within the domain of interest but the lack of dataset for such evaluation cases makes it difficult to take fundamental decisions for the application uses, such as the choice of an embedding model or whether to use English as a pivot language to perform the task.
We propose a new simple and efficient dataset generation and evaluation scenario that simulates real-world challenges of cross-lingual semantic retrieval in large text collections. Our approach involves the use of LLMs to create adversarial examples from parallel sentences that challenge the cross-lingual embedding capabilities of multilingual models. Concretely speaking, four distractor sentences are generated for each parallel target sentence. We require these sentences to be very similar to the original target sentence in terms of syntactic structure and surface word forms, but to be semantically dissimilar. Our proposed task, Cross-Lingual Semantic Discrimination (CLSD), involves identifying the correct target sentence from among four distractor sentences in the target language, given an original sentence in the source language.
To better understand how the introduced changes in the distractors influence the semantic similarities cross- and monolingually, we perform a fine-grained analysis and correlate the semantic changes with linguistic properties at the level of part-of-speech tags.
Contributions
1. Proposing a new task that emulates cross-lingual semantic search within large text collections, requiring only a set of parallel sentences for dataset creation. 2. Publishing four such datasets within the news domain for the language pair German-French. 3. Demonstrate that within the task, certain embedding models perform better in direct cross-lingual retrieval, while others benefit from using English as a pivot language. 4. Providing insights into the cross- and monolingual semantic representation of multilingual models by a linguistically informed analysis.
2 Related Work
| Original Sentence | Distractor Sentence 1 | Distractor Sentence 2 |
Die Linkspartei beschließt in Bonn ihr Programm zur Europawahl.
(The Left Party adopts its program for the European elections in Bonn.) |
Die Linkspartei beschließt in Bonn ihr Programm zur Bundestagswahl.
(The Left Party adopts its program for the Bundestag elections in Bonn.) |
Die Linkspartei verweigert in Bonn ihr Programm zur Europawahl.
(The Left Party refuses its program for the European elections in Bonn.) |
Là-bas, la perspective de la fin de la guerre reste toujours très éloignée.
(Over there, the prospect of the end of the war is still a long way off.) |
Là-bas, la fin de la guerre semble toujours très incertaine.
(Over there, the end of the war still seems very uncertain.) |
Là-bas, la perspective de la paix semble toujours très éloignée.
(Over there, the prospect of peace still seems a long way off.) |
Cross-Lingual Semantic Search
Cross-lingual semantic search is an umbrella term for tasks that involve searching for texts with similar or relevant semantics in another language. Examples of such cross-lingual tasks are information retrieval (CLIR) (Lawrie et al., 2023, 2024), question answering (CLQA) (Lewis et al., 2020), semantic text similarity (X-STS) (Cer et al., 2017), and bitext mining (Research, ; Zweigenbaum et al., 2017). As shown in MTEB (Muennighoff et al., 2023), multilingual embedding models vary in their performance for different cross-lingual semantic search tasks and language pairs. When selecting models for a use case, we often rely on results from other domains or language pairs, as direct evaluation results are not always available.
Distractor Generation
PAWS-X Yang et al. (2019); Zhang et al. (2019) generated adversarial examples by rule-based and backtranslation methods to measure the limitations of paraphrase identification models. More recently, XSim++ Chen et al. (2023) used rule-based text augmentation to generate synthetic examples for tuning LASER models (replacements by antonyms, swapping entities). Within semantic search, InPars (Bonifacio et al., 2022) use solely in-domain synthetic data to finetune retrieval models that surpass strong baselines.
| Model (Hugging Face Name) | X-Ling. Aligned | Paraph. Aligned | Retrieval Aligned | WMT19-DE->FR | WMT21-DE->FR | WMT19-FR->DE | WMT21-FR->DE | Mean |
| Cross-Lingual CLSD Evaluation | ||||||||
| multilingual-e5-base | ✓ | ✓ | ✓ | 91.51 | 86.34 | 88.46 | 81.97 | 87.07 |
| multilingual-e5-large | ✓ | ✓ | ✓ | 94.43 | 91.38 | 91.45 | 87.57 | 91.21 |
| gte-multilingual-base | ✓ | ✓ | ✓ | 90.22 | 90.48 | 89.55 | 91.60 | 90.46 |
| paraphrase-multilingual-mpnet-base | ✓ | ✓ | ✗ | 91.31 | 91.15 | 91.11 | 92.95 | 91.63 |
| sentence-transformers/LaBSE | ✓ | ✓ | ✗ | 95.18 | 94.06 | 94.30 | 94.18 | 94.43 |
| Machine Translation (M2M 1.2B Fan et al. (2021)) to EN -> CLSD Evaluation | ||||||||
| multilingual-e5-base | ✓ | ✓ | ✓ | 90.50 | 89.81 | 89.55 | 93.39 | 90.81 |
| multilingual-e5-large | ✓ | ✓ | ✓ | 90.22 | 91.83 | 89.95 | 92.95 | 91.24 |
| gte-multilingual-base | ✓ | ✓ | ✓ | 89.68 | 90.26 | 89.07 | 92.95 | 90.49 |
| paraphrase-multilingual-mpnet-base | ✓ | ✓ | ✗ | 87.92 | 88.13 | 87.64 | 91.15 | 88.71 |
| sentence-transformers/LaBSE | ✓ | ✓ | ✗ | 90.84 | 91.83 | 90.16 | 94.40 | 91.81 |
3 Experiments
3.1 Cross-Lingual Semantic Discrimination (CLSD)
We propose a task that aims to evaluate multilingual embedding methods based on their capacity to find the closest cross-lingual semantic match amidst large bilingual text collections. The goal is to measure the model’s ability to identify a parallel sentence as the most similar to the original sentence in the source language. To accomplish this task, a dataset of parallel sentences is required that will be enriched with distractor sentences in the target language.
In this paper, we test the hypothesis that LLMs can generate difficult examples that look similar on the surface, but convey different meanings. We prompt GPT-4 (gpt-4-0613) OpenAI (2023) to monolingually generate four distractors for each original sentence (see Figure 1). The distractors are requested to be structurally and lexically similar, whilst being semantically dissimilar. For each language, we manually examine 200 distractors of the original sentences and find that over 98% of the distractors meet these criteria. Example sentences are shown in Table 1.
CLSD Datasets
We generate and validate (samples of) four datasets, consisting of both language directions of German/French for the WMT19 Barrault et al. (2019) and WMT21 Akhbardeh et al. (2021) test sets.111The data is available under ANONYMURL.. In the 18,928 generated distractors, we observe that the LLM used different modification strategies such as single polarity inversion, entity replacement or even multiple swaps with word-level Jaccard similarity between originals and distractors at (, ). Table 4 illustrates samples of distractors sets.
CLSD Evaluation
To evaluate a multilingual embedding model , we assess its ability to discriminate between original parallel sentence pairs and challenging distractors. For a given sentence in the source language and its corresponding translation in the target language, along with a set of distractor sentences in the target language, the model must produce an embedding of that is more similar to the embedding of than to that of any distractor in .
To measure performance of embedding models, we use Recall@1 (R@1), defined as the proportion of samples in which the model ranks the true translation as more similar to the source sentence than the four distractors. For example, in Table 2, the column WMT19-DE->FR reports the R@1 of identifying the original French target sentence as more similar than its French distractors.
Models Evaluated
We examine the following five embedding models available through SentenceTransformers Reimers and Gurevych (2019).
multilingual-e5-base/-large (M-E5-b/l) Wang et al. (2024, 2024) transform XLM-RoBERTa Conneau et al. (2020) into a bi-encoder through weakly supervised contrastive pre-training with multilingual data and a second step of supervised fine-tuning on retrieval datasets.
gte-multilingual-base (M-GTE) Zhang et al. (2024) pre-train a multilingual long-context encoder (8192) and derive a bi-encoder by a process similar to multilingual-e5 that also includes contrastive training against hard negatives.
paraphrase-multilingual-mpnet-base-v2 (M-MPNet) Reimers and Gurevych (2020) use knowledge distillation through cosine loss training on parallel sentences between an English paraphrase trained teacher model (paraphrase-mpnet-base-v2) and a multilingual student model (XLM-RoBERTa Conneau et al. (2020)) to teach the student to represent sentences closely across multiple languages.
LaBSE Language-agnostic BERT Sentence Embedding Feng et al. (2022) is trained with translation ranking loss and negative samples.
3.2 CLSD Results
The results of the models on the CLSD datasets are shown in Table 2. Regarding the direct cross-lingual evaluation, the Bitext Mining specialist LaBSE is the best model in this task with an average score of 94.43 (R@1), demonstrating its consistent ability to cross-lingually rank exact parallel text higher than the distractors. Another pattern that emerges is that models (M-E5-b/l, M-GTE) that were also trained on English-centric retrieval datasets, such as the MS Macro Nguyen et al. (2016), perform worse within our stricter task.
In the lower half of the table, we evaluate the models using English as a pivot language on both sides, employing the M2M 1.2B translation model Fan et al. (2021). Both LaBSE and M-MPNet show a performance decrease of 2.6/2.8 points compared to direct cross-lingual evaluation. In contrast, models with weaker performance in direct cross-lingual show only minor improvements, except the M-E5-b model, which gains 3.7 points.
We conduct a qualitative analysis of the 279 samples where LaBSE succeeds in the direct cross-lingual evaluation but fails in the pivot-through-English evaluation. Our analysis reveals that, in most cases, language-specific terms are translated into broader and common English words, reducing the distinction between the original texts and the distractors. Less frequently, we observe omitted details in the translations, hallucinations, and, rarely, poor quality translations.
3.3 Fine-grained similarity analysis
How do minor surface changes affect the semantic similarity between sentence pairs? Can linguistic analysis provide a more precise characterisation of the changes? To address these questions, we pick original-distractor pairs in which the distractor differs from the original by only a single word. This subset includes 523 pairs (5.5%) in German and 324 pairs (3.4%) in French. Using Stanza Qi et al. (2020) and further validation by a linguist annotator, we assign part-of-speech tags to the differing token. We measure the cross-lingual semantic change by subtracting the cosine distance between the original source sentence and the target distractor from the cosine distance between the original parallel pair. To improve comparability in our analysis across different models, we normalize222The average cosine similarity difference between parallel and unrelated cross-lingual sentence pairs in the original WMT19/21 datasets is the normalization factor in Figure 2. this score by dividing with for the corresponding model. Formally, we measure:
Figure 2 presents model results, showing that the M-MPNet model experiences the largest overall shifts in cosine similarity when single-word modifications are made. The column “ANY” reflects changes across all items, facilitating within-model comparisons to determine which linguistically characterized subgroups exhibit more pronounced changes. Notably, LaBSE, the best performing model within CLSD, represents these single-word changes with relatively uniform similarity shifts across linguistic categories, unlike the other models. Additionally, we observe that replacing proper nouns consistently decreases semantic similarity more than replacing other parts of speech across all languages and models.
Mono- and Cross-Lingual Change Correlations
How language-agnostic are the base models when dealing with distractors? Does it matter whether the source sentence is presented in the source or target language? We measure the monolingual similarity change of the original to the distractor by substituting with in the calculation. The corresponding monolingual semantic change plots can be found in Figure 3 in the Appendix. By correlating the results of the cross- and monolingual change for all one-word change cases (i.e. “ANY”): The correlation between monolingual and cross-lingual changes is strong for M-MPNet with 0.93 closely followed by M-GTE with 0.85 and LaBSE with 0.82. The M-E5-b/l shows lower correlation with 0.79 and 0.71 respectively. However, specific nuances depend on the language, model, or type of replaced word. For example, M-GTE shows a low correlation of 0.64 for proper nouns between cross- and monolingual evaluations in French, compared to 0.77 for German.
4 Conclusion
This paper contributes to cross-lingual semantic search by identifying gaps in current evaluation practices and introducing adversarial examples to test multilingual embeddings. Our research, on the German-French case study, introduces four new adversarial cross-lingual evaluation datasets for the news domain. Our comparative evaluation of direct cross-lingual and pivot through English evaluation highlights the advantage of each method based on the embedding model used. The fine-grained analyses demonstrate how generated distractors offer insights into the cross-lingual capabilities of multilingual models and detail how specific small perturbations in sentences can change these models’ produced semantic representations. Future work should extend these approaches to other language pairs and refine adversarial techniques to improve semantic search model evaluations.
Limitations
We introduce a pipeline that allows cross-lingual evaluation of semantic search between any two languages that have an available set of parallel texts. However, we evaluate only a single language pair (of interest) and only a single model for distractor generation, whereas in reality much cheaper methods might suffice. Automatic prediction of PoS tags and named entities in our analysis can result in introducing noise within the fine-grained analysis results. The performance of the stanza models and our added manual results validation minimises this risk. However, the capabilities of current models in the language of interest should be critically assessed. Furthermore, it is important to note that the examination of only the few models is due to space constraints; however, it is worth considering that valuable insights can also be gained from smaller models. We hope that larger future work could address these limitations and explore this approach more broadly, creating a large scale cross-lingual semantic search dataset parallel in many, including low-resource languages.
Acknowledgments
This research is conducted under the project Impresso – Media Monitoring of the Past II Beyond Borders: Connecting Historical Newspapers and Radio. Impresso is a research project funded by the Swiss National Science Foundation (SNSF 213585) and the Luxembourg National Research Fund (17498891).
Appendix A Appendix
| Model Parameter | Value |
| Temperature | 1.0 |
| Top-P | 1.0 |
| Distractor | Intra-Distractor Jaccard Sim. |
| Original: PARIS (Reuters) - Die New Yorker Börse schloss am Freitag mit 0,68 % im Minus und beendete die Woche mit einem allgemeinen Rückgang, da sie von enttäuschenden Ergebnissen, einem Wiederanstieg der Coronavirus-Infektionen und geopolitischen Unsicherheiten beeinträchtigt wurde. | / |
| Adv1: LONDON (Reuters) - Die Londoner Börse schloss am Dienstag mit 0,68 % im Plus und startete die Woche mit allgemeinem Wachstum, da sie von erfreulichen Ergebnissen, einem kontinuierlichen Rückgang der Coronavirus-Infektionen und geopolitischen Stabilitäten angetrieben wurde. | 0.434 |
| Adv2: BERLIN (Reuters) - Die Berliner Börse schloss am Donnerstag mit 0,68 % im Minus und beendete die Woche mit einem insgesamt flachen Verlauf, das durch gemischte Ergebnisse, eine Stagnation der Coronavirus-Infektionen und geopolitische Spannungen beeinflusst wurde. | 0.300 |
| Adv3: TOKYO (Reuters) - Die Tokioter Börse schloss am Mittwoch mit 0,68 % im Plus und begann den Monat mit allgemeiner Erholung, da sie von positiven Ergebnissen, einem eutlichen Rückgang der Coronavirus-Infektionen und geopolitischer Sicherheit profitierte. | 0.359 |
| Adv4: MADRID (Reuters) - Die Madrider Börse schloss am Montag mit 0,68 % im Minus und startete die Woche mit allgemeinem Rückgang, da sie von enttäuschenden Ergebnissen, einem Wiederanstieg der Coronavirus-Infektionen und geopolitischen Unsicherheiten beeinträchtigt wurde. | 0.438 |
| Original: Der Nasdaq verzeichnete die schlechteste Woche der letzten vier. | / |
| Adv1: Der Nasdaq hat die beste Woche der letzten vier verzeichnet. | 0.630 |
| Adv2: Der Nasdaq verzeichnete die aktivste Woche der letzten vier. | 0.623 |
| Adv3: Der Nasdaq hat die ruhigste Woche der letzten vier verzeichnet. | 0.630 |
| Adv4: Der Nasdaq verzeichnete die turbulenteste Woche der letzten vier. | 0.623 |
| Original: Die Beamten werden im Haushalt für 2021 jedoch nicht vergessen werden. | / |
| Adv1: Die Beamten werden im Haushalt für 2021 jedoch nicht besprochen werden. | 0.818 |
| Adv2: Die Beamten werden im Haushalt für 2021 jedoch nicht entlassen werden. | 0.818 |
| Adv3: Die Beamten werden im Haushalt für 2021 jedoch nicht befördert werden. | 0.818 |
| Adv4: Die Beamten werden im Haushalt für 2021 jedoch nicht belastet werden. | 0.818 |
| Original | Distractor | Tricked Model |
| WMT19- DE->FR-CLSD | ||
| Kipping au congrès de die Linke sur l’Europe : l’Europe est depuis longtemps un continent d’immigration. | Kipping au congrès de die Linke sur l’Asie : l’Asie est depuis longtemps un continent de diversité. | MPNet-M |
| Par exemple, nous allons à présent recevoir un quartier général pour les munitions de l’UE, à Bruxelles, ce que les Britanniques ont jusqu’ici toujours empêché. | Par exemple, nous prévoyons de lancer un bureau principal pour les ressources de l’UE, à Bruxelles, ce que les Français ont toujours refusé jusqu’ici. | MPNet-M |
| L’appel de Macron: l’Europe est davantage qu’un "projet" | L’appel de Macron: l’Europe est plus qu’une "idée". | M-E5-b |
| Trois commissions ont accepté ma candidature, l’une d’elle était de justesse contre. | Trois comités ont rejeté ma candidature, l’un d’eux était presque en faveur. | M-E5-b |
| WMT19- FR->DE-CLSD | ||
| Würden sie dies tun, gäbe es niemanden in Europa, der sich dagegen wehren würde. | Würden sie dies tun, gäbe es niemanden in Europa, der sich dafür aussprechen würde. | MPNet-M M-E5-b |
| Um dem etwas entgegenzusetzen, ist die EU-Kommission auf Kooperation mit den Internetriesen angewiesen. | Um das zu fördern, ist die EU-Kommission auf Zusammenarbeit mit den Internetgiganten angewiesen. | MPNet-M M-E5-b |
| Und in der Sozialpolitik sehe ich bei ihr ebenfalls keinen Rechtsschwenk. | Und bei ihr sehe ich in der Sozialpolitik ebenfalls keinen Linksschwenk. | MPNet-M |
| Bohnentopf, Hähnchen mit Pommes oder Kalbfleisch mit Soße und Kartoffeln: In der Mensa-Eingangshalle der Universität Karlsruhe drängeln sich um die Mittagszeit die jungen Leute. | Pizzaschnitten, Hähnchen mit Reis oder Rinderbraten mit Gemüse und Kartoffeln: Um die Abendstunde füllt sich der Festsaal der Universität Stuttgart mit hungrigen Studenten. | MPNet-M |
| Nein. Zurückweisungen gibt es ja auch in anderen EU-Ländern. | Nein, Zuwanderungen gibt es ja auch in anderen EU-Ländern. | M-E5-b |
| Und es ist zugleich die Emanzipation der Linken von ihrem Ex-Vorsitzenden Oskar Lafontaine. | Und es ist zugleich die Spaltung der Linken wegen ihrem Ex-Vorsitzenden Oskar Lafontaine. | M-E5-b |